estimate the
roots of the equation: y = x^2 - 2x - 8

Answer by ewatrrr(24785)


Hi, previously posted.
Note: roots of an equation are found when y = 0
x^2 - 2x - 8 = 0
Factoring
(x-4)(x+2) = 0 Note:SUM of the inner product(-4x) and the outer product(2x) = -2x
(x-4) = 0 x = 4
(x+2) = 0 x = -2